OpenStudy (anonymous):

Shiming, Doreen, and Brian are studying independently for their driver’s license test. The probability Shiming will pass the test is 3/5, the probability Doreen will pass the test is 3/28, and the probability Brian will pass the test is 3/10. What is the probability at least one of these people passes the test?

ganeshie8 (ganeshie8):

i think we can find probability for NONE of them will pass. and subtract it from 1

ganeshie8 (ganeshie8):

probability for none of them passing = (1-3/5)(1-3/28)(1-3/10)

OpenStudy (anonymous):

So it's 2/5 * 25/28 * 7/10?

ganeshie8 (ganeshie8):

1- that

OpenStudy (anonymous):

So it's 7/8?

ganeshie8 (ganeshie8):

wolfram says 3/4

ganeshie8 (ganeshie8):

pls do check if there is some flaw in my logic.. . im feeling slight dyslexic today...

OpenStudy (anonymous):

Let's see...your logic is correct, but I don't understand it...

ganeshie8 (ganeshie8):

i thought like this : P(atleast one of them passing) + P(none of them passing) = 1

ganeshie8 (ganeshie8):

P(atleast one of them passing) = 1 - P(none of them passing)

ganeshie8 (ganeshie8):

P(none of them passing) = P(A') . P(B') . P(C') note that these are mutually exclusive evens. one guy passing doesnt depend on other guys passing...

OpenStudy (anonymous):

But why did you have me subtract one form A, B, and C and then subtract again?

ganeshie8 (ganeshie8):

becos, P(A) + P(A') = 1

ganeshie8 (ganeshie8):

we are given P(A) <---- p(A passing) we need P(A') <------ p(A not passing)

ganeshie8 (ganeshie8):

P(atleast one of them passing) = 1 - P(none of them passing) P(none of them passing) = P(A') . P(B') . P(C')

ganeshie8 (ganeshie8):

we have used both above equations

OpenStudy (anonymous):

Oh. ok...now I get it...first you subtract the pass rate from 1 to get the non pass rate individually. Then multiply them to get the combined no pass rate, then subtract from 1 again to get the combined pass rate?

ganeshie8 (ganeshie8):

thats exactly what we did :)

OpenStudy (anonymous):

ALright THank you :)

ganeshie8 (ganeshie8):

np.. :)